OpenStudy (anonymous):

Can I please get someone to CHECK over my work to see if I did it properly? I will still award medal.

OpenStudy (anonymous):

ya say

OpenStudy (anonymous):

3. Robbie the Robot is on a weather satellite orbiting Earth about 3600 km above the surface. The Earth’s radius is about 6400 km. He has had a malfunction in his output device, and the satellite is traveling without communication. His last report was only in terms of trigonometric values and was only partially received. It said, sin Θ < 0……and then he was lost again. Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning. (8,000, -6,000) is a reasonable coordinate because it sits on the circle with a radius of 10,000. This is the radius from the center of the earth (the source of gravity that controls the orbit) and the satellite because “Earth about 3600km above the surface.” and “The Earth’s radius is about 6400km.” And 3600+6400=10,000; hence the point (8,000, -6,000) is reasonable because it is sitting on the circumference of the orbit. Also, either the x-intercept or the y-intercept of the point chosen must be negative because sin Θ < 0, and sin is only negative in the 3rd and 4th quadrants of the graph. Part 2: What are the values of the sin Θ, cos Θ, and tan Θ using your coordinate point? The only information we need to find the sin Θ, cos Θ, and tan Θ is the following: Y = -6,000 X = 8,000 R = 10,000 Sin Θ = y/r Cos Θ = x/r Tan Θ = y/x Sin Θ = -6, 000/10,000 Cos Θ = 8,000/10,000 Tan Θ = -6,000/8,000 Sin Θ = -.06 Cos Θ = .08 Tan Θ = -0.75

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks. May I ask how you know this?

OpenStudy (anonymous):

i ll catch u later m bsy sry!

OpenStudy (anonymous):

OK. Let me re-phrase: I need a RELIABLE source to tell me I'm right.

OpenStudy (anonymous):

get lost

OpenStudy (anonymous):

Why are you still here? I thought you were too busy.

OpenStudy (anonymous):

nw listn u feather...\[\sqrt{3600+6400} is the hypotenuse and hence \sin \alpha is perpendicular/hypotenuse \in the same way \cos \alpha is base/hypotenuse \]

OpenStudy (anonymous):

Openstudy is not a place for big jerks to call people names. If you don't want to help me, go away.

OpenStudy (anonymous):

i wrote ur ans

OpenStudy (anonymous):

There are nice people who will help me.

OpenStudy (anonymous):

Ok. Since you can't tell me how you know that, I am not going to give you a medal.

OpenStudy (anonymous):

so did i write in spanish??

OpenStudy (anonymous):

Please go help someone who does not have feelings.

OpenStudy (anonymous):

ohk ohk..m sry k? u ask ur ques step by step i ll ans

OpenStudy (anonymous):

I already answered it. I was looking for more like..."Hey, GreenEyes, I am a math teacher...What you have done is right."

OpenStudy (anonymous):

den wat i remember is i said "yup" in the first cmmnt

OpenStudy (anonymous):

And I asked for confirmation that you knew what you were doing. How do you know that I am right?

OpenStudy (anonymous):

You could be just looking for an easy medal, not that I'm saying that you are.

OpenStudy (anonymous):

ur reasons r correct cuz it satisfies the conditions

OpenStudy (anonymous):

O, an' ur speling tels me that u r a reliable source. Thx fer ur effort, but idk if u realy no the info. I will just go with what I have because not getting a second opinion is less stressful to me than dealing with you. I will pray for you to have better, happier days in the future.

OpenStudy (anonymous):

So you think you are smarter...the sentence is less time consuming...so i prefer that language...better dont dare my english..and ya if you are that intelligent...then why dont you challenge me

OpenStudy (anonymous):

because everyone aint free like u..